OpenStudy (anonymous):

@jigglypuff314 can you help i dont understand this?

OpenStudy (anonymous):

With what?

OpenStudy (anonymous):

The table below shows some values of f(x) and g(x) for different values of x. x f(x)=2x+24 g(x)=5(0.5)^x -2 20 20 -1 22 10 0 24 5 1 26 2.5 2 28 1.5

OpenStudy (anonymous):

i got the graph part down i just dont understand the question

OpenStudy (anonymous):

Complete the chart and determine the solution of the equation f(x) = g(x)? x = -2 x = -1 x = 1 x = 20

OpenStudy (anonymous):

i think its 20

jigglypuff314 (jigglypuff314):

I'm not good with x as the exponent... but it would be to solve for x when 2x+24 = 5(0.5)^x

OpenStudy (anonymous):

i dont need you to solve that i already did it its the numbers underneath the equations i just dont understand what the questions asking me

jigglypuff314 (jigglypuff314):

it's just asking for x then right?

OpenStudy (anonymous):

idk thats what im comfused